Wildhorse Company purchased machinery on January 1, 2020, for $94,400. The machinery is estimated to have a salvage value ofPART B) Compute 2020 depreciation expense using the double-declining-balance method, assuming the machinery was purchased on October 1, 2020.
Answer #1

Double declining rate = 200/useful life = 200/8 = 25%

A)

Depreciation expense = 94,400*25%

= 23,600

B)

Depreciation expense = 94,400*25%*3/12

= 5900